OpenStudy (anonymous):

@AnswerMyQuestions @iGreen estimate the sum or diffrence. 4/5+ 18/20=

OpenStudy (igreen):

4/5 is 0.8 18/20 is 0.9 What's 0.8 + 0.9?

OpenStudy (anonymous):

0.8 + 0.9

OpenStudy (igreen):

Copies me ^^

OpenStudy (igreen):

Can you add 0.8 + 0.9? @msdf!!!!!

OpenStudy (anonymous):

1.7

OpenStudy (igreen):

Yep, so that's your answer.

OpenStudy (igreen):

Do you have options? 'Cause it does say to estimate..and 1.7 is the direct answer.

OpenStudy (anonymous):

i green the answer choice are A. 1 B.1/2 2

OpenStudy (igreen):

Okay, is 1.7 closer to 1 or 2?

OpenStudy (igreen):

or 1/2?

OpenStudy (anonymous):

1

OpenStudy (igreen):

Check again.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

1/2

OpenStudy (igreen):

2 - 1.7 = 0.3 1.7 - 1 = 0.7 1.7 - 0.5 = 1.2 Which one is smaller?

OpenStudy (igreen):

0.3, 0.7, or 1.2?

OpenStudy (anonymous):

0.3

OpenStudy (igreen):

Yep, so what's your answer going to be? 2 - 1.7 = 0.3

OpenStudy (anonymous):

2

OpenStudy (igreen):

Yep!